OpenStudy (anonymous):

\[\sqrt{2x^2 +9} - \sqrt{2x^2-9} = \sqrt{a} -\sqrt{b} , and \sqrt{2x^2+9} + \sqrt{2x^2-9} = 9 - 3\sqrt{7}\] then the values of a and b are respectively?

OpenStudy (anonymous):

error

OpenStudy (anonymous):

where

OpenStudy (asnaseer):

\[\sqrt{2x^2+9}+\sqrt{2x^2-9}=9-3\sqrt{7}\]\[(\sqrt{2x^2+9}+\sqrt{2x^2-9})(\sqrt{2x^2+9}-\sqrt{2x^2-9})=(9-3\sqrt{7})(\sqrt{2x^2+9}-\sqrt{2x^2-9})\]\[(2x^2+9)-(2x^2-9)=(9-3\sqrt{7})(\sqrt{2x^2+9}-\sqrt{2x^2-9})\]\[18=(9-3\sqrt{7})(\sqrt{2x^2+9}-\sqrt{2x^2-9})\]therefore:

OpenStudy (asnaseer):

\[\sqrt{2x^2+9}-\sqrt{2x^2-9}=\frac{18}{9-3\sqrt{7}}=\frac{18(9+3\sqrt{7})}{(9-3\sqrt{7})(9+3\sqrt{7}))}\]\[\qquad=\frac{18(9+3\sqrt{7})}{81-73}=\frac{18(9+3\sqrt{7})}{18}=9+3\sqrt{7}\]

OpenStudy (asnaseer):

so now we have two equations that we can use:\[\sqrt{2x^2+9}+\sqrt{2x^2-9}=9-3\sqrt{7}\]\[\sqrt{2x^2+9}-\sqrt{2x^2-9}=9+3\sqrt{7}\]

OpenStudy (anonymous):

the answer is 81, 63 @asnaseer did u got that

OpenStudy (asnaseer):

if we add these up we get:\[2\sqrt{2x^2+9}=18\]\[\sqrt{2x^2+9}=9\]\[2x^2+9=81\]therefore:\[2x^2-9=63\]therefore:\[\sqrt{2x^2+9}-\sqrt{2x^2-9}=\sqrt{81}-\sqrt{63}\]and you have your answer.

OpenStudy (anonymous):

let me check and if i have doubts i will inform u thanzzzzz

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

i got it now thzz

OpenStudy (asnaseer):

ok - good :)

OpenStudy (asnaseer):

I've been checking this result myself as it seemed odd that: \[\sqrt{2x^2+9}+\sqrt{2x^2-9}=9-3\sqrt{7}\]\[\sqrt{2x^2+9}-\sqrt{2x^2-9}=9+3\sqrt{7}\] I think your second equation should have been:\[\sqrt{2x^2+9} + \sqrt{2x^2-9} = 9 + 3\sqrt{7}\]

OpenStudy (anonymous):

ya it 2 +

OpenStudy (anonymous):

u wrote 2*

OpenStudy (asnaseer):

i.e. the question should have been: \[\sqrt{2x^2 +9} - \sqrt{2x^2-9} = \sqrt{a} -\sqrt{b} , and \sqrt{2x^2+9} + \sqrt{2x^2-9} = 9 + 3\sqrt{7}\]

OpenStudy (anonymous):

no it is 9-3root7..